OpenStudy (anonymous):

calculate the number of C, H and O atoms in 1.50 g of glucose (C6H12O6) a sugar.

OpenStudy (somy):

can u find mole of C6H12O6 ?

OpenStudy (anonymous):

is it 8.3 x 10^-3 mole

OpenStudy (somy):

mole= mass/Mr

OpenStudy (somy):

yes you are right

OpenStudy (abhisar):

1 mole of C6H12O6 will contain 6 moles of C, 12 moles of Hydrogen & 6 moles of oxygen. M i correct @Somy ?

OpenStudy (abhisar):

Now all u hv to do is to find no of moles in 1.50 g C6H12O6

OpenStudy (somy):

yes

OpenStudy (somy):

but its not mole its atom

OpenStudy (abhisar):

Hmm..we can find number of atoms from moles by multiplying it by avagadro number

OpenStudy (somy):

no not yet

OpenStudy (somy):

now what you will to is this (Mr of C alone * mole that you just got)/ total Mr

OpenStudy (somy):

and you will get mole of C

OpenStudy (somy):

do make sure to take Mr of all 6C atoms meaning 12*6

OpenStudy (abhisar):

8.33 * 10\(^{-3}\) moles of glucose will produce 6*8.33 * 10\(^{-3}\) moles of carbon

OpenStudy (somy):

ummm no

OpenStudy (somy):

12*6= 72 is Mr of C alone Mr of the whole compound is 180 and mole is 8.3 x 10^-3 so mole of C= (72* 8.3*10^-3)/ 180 =

OpenStudy (somy):

do same thing with O and H and you will get moles for all of them then multiply each mole for each element by 6.02*10^23 which is Avogadro number

OpenStudy (abhisar):

1 mole of water will contain how many moles of hydrogen ?

OpenStudy (somy):

H2O if u say is 1 mole then mole of H would be 2*mole/Mr(H2O)

OpenStudy (abhisar):

Then how was i wrong ?

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar 8.33 * 10\(^{-3}\) moles of glucose will produce 6*8.33 * 10\(^{-3}\) moles of carbon \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

1 mol of water contains 6.022*10^23 molecules. Each molecule of water contains 2 atoms of hydrogen: Therefore there are 2*6.022*10^23 atoms = 1.2044*10^24 atoms of hydrogen in 1 mol of water.

OpenStudy (abhisar):

correct ?

OpenStudy (somy):

~_~ AB either you are not understanding it or i am doing it wrong @

OpenStudy (somy):

@chmvijay

OpenStudy (abhisar):

Its glucose not sugar(C12H22O11)

OpenStudy (abhisar):

ohkay lets see, btw wt do u mean by Mr ?

OpenStudy (somy):

molecular mass

OpenStudy (chmvijay):

@Biha did u get it what they are saying so far?

OpenStudy (somy):

@chmvijay waaaaaaaaait =_= man i asked you if im doing wrong and he is doing right or opposite way round @Abhisar is my 2tag teammate ♥

OpenStudy (somy):

and thnx btw @chmvijay ♥